INAF queue observing, 2015 Jun 22-23

Observer: F. Cusano, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Olga Kuhn on call
Instrument: MODS, LUCI

Summary

At the sunset some cloud at the horizon, humidity 30%
At 6:21 the TO closes the dome for clouds and chance of rain.
The TO opens the dome at 07:04, some cirrus up to 10 then clear sky.

Night duration (12\xB0 el twilight) 20:30-4:00 ( 3:30-11:00 UT) = 7h:30m

Observing time: 5 hrs

Weather time loss: 2.0 hrs (27%) Techical problem: 0.3 hrs (4%)

Data Summary

At the sunset some cloud at the horizon, humidity 30%

Project_RUN Objects Filter ExpTime Notes
Standard spec Hz44 dual grating, red and blue grating    
Sommariva ngc 4579 dual grating 23 m clouds, check
Lardo M92 blue grating 22 m stop for clouds, but good data
Israel A5 cxouj215447 dual grating 2h completed 10 min with clouds
Piconcelli B3 2123-0050 210_zjHK 15m completed
telluric B3   210_zjHK    

Details

03:10 We slew to MODS standard Hz44, Della Ceca B program
  • acqMODS Standard_11B.acq mods1r.20150623.0002-0003
  • 03:27 execMODS Standard_11B.obs,
  • dual grating mods1r.20150623.0004-0006; mods1b.20150623.0002-0004
  • red grating mods1r.20150623.0007-0009, seeing 1.1, cirrus coming in all over the sky, guide star fainting
  • blue grating mods1b.20150623.0005-0007

04:03 weather conditions are not optimal for the faint target Della CecaB, cirrus cover all the sky, seeing 1.2

04:04 we slew to Sommariva NGC4579
  • acqMODS ngc4579_mods.acq mods1r.20150623.0010-0013, problem with acquistion too mouch clouds, problem solved

04:28 the sky is completely covered by clouds

04:48 conditions are better, we send ngc4579_mods.obs
  • 05:03 stop integration too much clouds mods1r.20150623.0014; mods1b.20150623.0008 not good
  • 05:08 send integration again, guide star again visibile in the guider
  • 05:18 the sky is almost clear, seeing 1.1
  • 05:32 stop integration object too low on the horizon, texp=1350s , mods1r.20150623.0015; mods1b.20150623.0009 are good

05:35 We slew to Lardo M92
  • acqMODS mods.16.M92.acq mods1b.20150623.0010-0012
  • seeing 1.1

05:49 execMODs mods.16.M92.obs, mods1b.20150623.0013, good
  • seeing 1.0
  • clouds during the first integration
  • 06:17 lost guide star, stop exposure texp=1300, spectra look good

06:21 The TO closes the dome

07:04 The TO opens the dome, the sky is 80% clear

07:15 we slew to Israel A5
  • acqMODs catsj215447_ls.acq, object very faint we increse the exposure time of the acquisition two times
  • mods1r.20150623.0016-0020

07:47 execMODS catsj215447_ls.obs, some thin cirrus
  • mods1r.20150623.0021-0023; mods1b.20150623.0014-0016
  • seeing 1.0
  • note we place the object at y=628 in the acquisition and in 2d spectra frame is at y=1649
  • some clouds from 9:05

09:56 swith to LUCI

10:19 TO solved problem with calibration unit locked inside

10:25 executeLUCI..sh B3_LS.acq, Piconcelli B3
  • luci.20150623.0004-0007
  • seeing 0.8

10:45 executeLUCI..sh B3_ls.spec, Piconcelli B3
  • luci.20150623.008-0010

11:05 acquistion standard B3, tellB3_bis_ls_jHeI.acq
  • luci.20150623.0011-0014

11:10 execut... tellB3_ls_210_H.spec
  • luci.20150623.0015-0016

11:28 calibration with MODS
  • ISRAEL A5 flats mods1r.20150623.0024-0028; mods1b.20150623.0017-0021
  • ISRAEL A5 arc mods1r.20150623.0029, mods1b.20150623.0022
  • Lardo M92 file calibration for the mask mods1b.20150623.0023-0027
  • Sommariva file calibration for the mask ngc4579 mods1r.20150623.0028-0034; mods1b.20150623.0030-0032
  • red grating calibration, slitless flat grpixflats_r.cal mods1r.20150623.0035-0039
  • slitflats_5as_r.cal, flat slit 5 arcsec red grating, mods1r.20150623.0040-0042
  • grlamps_r.cal, lamps for red grating, mods1r.20150623.0043-0045

-- %USERSIG{RobertaCarini - 2015-06-23}%

Comments

 
Topic revision: r5 - 25 Jun 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback